# Graph Topology Estimation of Power Grids using Pairwise Mutual Information of Time Series Data

### v2

@@ 3.4.2 Rounding Data to a Significant Digit
Besides discretizing our data to a certain amount of bits we can also round our data to a significant digits of voltage magnitude data. Reducing the significant digits required will allow smart-meter designers and allow utility companies to reduce the size of data that need to be stored. The data is rounded the most significant digit of analysis prior to to discretizing the data. This method has similarities to the number of bits analysis, however, the number of bits analysis rounds the continuous data to the closest integer bit value after scaling, whereas the following analysis simply rounds to the closest integer without scaling. In Figure 14, we can see the effect of using SG1 data to the nearest significant digit on the SDR. Note that the Gaussian method is rounded before analysis to a significant digit. The other MI methods have data multiplied by a base ten number and then rounded to nearest integer before discrete MI method analysis (this is done since the JVHW and discrete MI methods only accept positive integers as input due to their implementation in this work). The Gaussian method surprisingly outperforms the other MI methods. For SG1, it’s clear that only millivolts (mV) of data are needed for all MI methods. The Gaussian method outperforms the other method needing 10’s of mV of accuracy. For SG1 (solar), the leaf SDR shows all methods need only 10’s of millivolts of data to perform optimally.
@@
In Figure 15, we see for SG2 (non-solar), the Gaussian methods outperform other MI methods at all precision levels except 0 V. For SG2 (solar), the Gaussian method performs similarly to to other methods and slightly better when less precision of data is used. The leaf SDR is affected significantly by the number of significant digits. For SG2, the Gaussian performs the best. For SG2 (solar), the Gaussian method performs similarly to the other methods, except it outperforms the other methods when using precision of 10 micro-Volts of data and 100’s of millivolts of data. We might expect the discrete MI methods to outperform the Gaussian method since the Gaussian approximation can be a crude approximation of the data. Both the SG2 and SG1 solar and non-solar results suggest precision of at least millivolts is required for estimation resulting in above $90\%$ SDR.
@@ 4 Discussion
This work performs experiments on different data sets, different lengths of data, different down-sampling schemes, different levels of precision of data, and different methods to caclualte the mutual information to visualize and quantify the performance of the Chow-Liu algorithm. The algorithm predicts the grid 100% correctly for most IEEE data sets and for the SG1 data sets (GridLAB-D). The algorithm however finds more difficulty with the larger 200+ node network in SG2 data sets (GridLAB-D) with greater than 90% of branches in the true system correctly estimated. The drop in performance for larger networks suggests the need for more experiments with SG2-data sets and other similar size data sets.
@@
Future work should focus on the further error frequency analysis for SG2 data sets, more use of a mixed Gaussian mutual information based algorithm, creating improved discretization methods. Lastly, state-of-the-art machine learning methods such as graph neural networks Wu et al. (2020), transformers Mialon et al. (2021), and neural architecture searches Zoph and Le (2016); Speckhard et al. (2023) to find the best architectures have found use in a wide variety of uses in the power grid modeling field Varbella et al. (2024); Mabrouk and Rajagopal (2022) and may be applicable to these datasets and should be investigated in future work.
@@
The code was then written in python. The repository can be found at https://github.com/speckhard/python_grid_est. An 8 node IEEE example is shown in an iPython notebook file in the example folder as well as an example main file. The repository can be installed using pip https://pypi.org/project/grid_top_est/.
@@ Appendix A Additional Plots and Data
The pairwise mutual information of the change in the current phasor for the SG1 data set in seen in Figure 16, The behavior appears to follow an exponential decay model.
@@
Figure 18: The true graph of the SG2 data set once redundant nodes are removed. This is the graph the algorithm will estimate and with which its performance will be compared. Note, this graph has a special layout and the text labels are removed since the graph has so many nodes that is difficult to fit in labels and use a non-compact layout.
@@
Node 23’s mutual information rankings in the SG1 dataset for the Gaussian approximations for different downsampling resolutions can be visualized in Figure 23 shows Node
@@
Figure 24 shows Node 23 for the SG1 solar datset using the Gaussian approximation for different downsampling resolutions.
@@
The different time windows effect the SDR for the SG1 (non-solar) dataset are shown in Figure 26.
@@
The histogram and Gaussian approximation of the change in voltage magnitude for Node 15 of the SG2 solar dataset is shown in Figure LABEL:SG2_solar_node15_deriv_hist. This node is consistently predicted to have an incorrect branch by the algorithm when using the Gaussian MI method for varying data lengths. This may be due to the Gaussian approximation being a poor one as seen in the Figure.
@@
The leaf SDR is analyzed is shown for 120 day time windows for the SG1 dataset in Figure 34.
@@ Appendix B Kruskal’s Algorithm
Kruskal’s algorithm seeks to find the minimum weight spanning tree. A tree is a graph with no cycles. In a tree you cannot draw a path using connections between nodes on the graph from one node and back to the same node node. Each connection is referred to as a pair of nodes $(x,y)$ in the graph and has an associated numerical value known as the weight $w(x,y)$ Cormen et al. (2022). The minimum weight spanning tree is the graph which connects the nodes so that the sum of the weights of the connections is minimized so long as there are no cycles in the graph.
@@
There are several different implementations of Kruskal’s algorithm, two of which differ in how they store the parent of a node Kershenbaum and Van Slyke (1972). When a pair of nodes is connected in the algorithm, one node is assigned as the parent of the other. This means one node has a higher ranking than the other. A node with four parents could be referred to as a rank four node. Alternatively, if a node has more than one parent, the path to the lowest ranked node can be compressed where the algorithm only keeps track of the lowest ranked parent for each node.
@@
Matlab has it’s own min-span-tree function but it requires the bio-informatics toolbox which many labs do not have. In order to increase usage of the grid topology estimation algorithm we wrote our own script to avoid licensing issues. Our implementation performs 10x slower than Matlab’s version as seen in Table 4. Since Kruskal’s algorithm execution time is negligible compared to mutual information computation time (as will be soon shown), minimal effort has been placed to improve our implementation of Kruskal’s algorithm.

### v3

@@ 3.4.2 Rounding Data to a Significant Digit
Besides discretizing our data to a certain amount of bits we can also round our data to a significant digit of voltage magnitude data. Reducing the significant digits required will allow smart-meter designers and allow utility companies to reduce the size of data that need to be stored. The data is rounded to the most significant digit of analysis prior to discretizing the data. This method has similarities to the number of bits analysis, however, the number of bits analysis rounds the continuous data to the closest integer bit value after scaling, whereas the following analysis simply rounds to the closest integer without scaling. In Figure 14, we can see the effect of using SG1 data to the nearest significant digit on the SDR. Note that the Gaussian method is rounded before analysis to a significant digit. The other MI methods have data multiplied by a base ten number and then rounded to nearest integer before discrete MI method analysis (this is done since the JVHW and discrete MI methods only accept positive integers as input due to their implementation in this work). The Gaussian method surprisingly outperforms the other MI methods. For SG1, it’s clear that only millivolts (mV) of data are needed for all MI methods. The Gaussian method outperforms the other method needing 10’s of mV of accuracy. For SG1 (solar), the leaf SDR shows all methods need only 10’s of millivolts of data to perform optimally.
@@
In Figure 15, we see for SG2 (non-solar), the Gaussian methods outperform other MI methods at all precision levels except 0 V. For SG2 (solar), the Gaussian method performs similarly to other methods and slightly better when less precision of data is used. The leaf SDR is affected significantly by the number of significant digits. For SG2, the Gaussian performs the best. For SG2 (solar), the Gaussian method performs similarly to the other methods, except it outperforms the other methods when using precision of 10 micro-Volts of data and 100’s of millivolts of data. We might expect the discrete MI methods to outperform the Gaussian method since the Gaussian approximation can be a crude approximation of the data. Both the SG2 and SG1 solar and non-solar results suggest precision of at least millivolts is required for estimation resulting in above $90\%$ SDR.
@@ 4 Discussion
This work performs experiments on different datasets, different lengths of data, different down-sampling schemes, different levels of precision of data, and different methods to calculate the mutual information to visualize and quantify the performance of the Chow-Liu algorithm. The algorithm predicts the grid 100% correctly for most IEEE datasets and for the SG1 datasets (GridLAB-D). The algorithm however finds more difficulty with the larger 200+ node network in SG2 datasets (GridLAB-D) with greater than 90% of branches in the true system correctly estimated. The drop in performance for larger networks suggests the need for more experiments with SG2-datasets and other similar size datasets.
@@
Future work should focus on the further error frequency analysis for SG2 datasets, more use of a mixed Gaussian mutual information based algorithm, creating improved discretization methods. Lastly, state-of-the-art machine learning methods such as graph neural networks [wu2020comprehensive], transformers [mialon2021graphit], and neural architecture searches [zoph2016neural, speckhard2023neural] to find the best architectures have found use in a wide variety of uses in the power grid modeling field [varbella2024powergraph, mabrouk2022distribution] and may be applicable to these datasets and should be investigated in future work.
@@
The code was then written in Python. The repository can be found at https://github.com/speckhard/python_grid_est. An 8 node IEEE example is shown in a Jupyter notebook file in the example folder as well as an example main file. The repository can be installed using pip: https://pypi.org/project/grid_top_est/.
@@ Appendix A Additional Plots and Data
The pairwise mutual information of the change in the current phasor for the SG1 dataset is seen in Figure 16, The behavior appears to follow an exponential decay model.
@@
Figure 18: The true graph of the SG2 dataset once redundant nodes are removed. This is the graph the algorithm will estimate and with which its performance will be compared. Note, this graph has a special layout and the text labels are removed since the graph has so many nodes that it is difficult to fit in labels and use a non-compact layout.
@@
Node 23’s mutual information rankings in the SG1 dataset for the Gaussian approximations for different downsampling resolutions can be visualized in Figure 23.
@@
Figure 24 shows Node 23 for the SG1 solar dataset using the Gaussian approximation for different downsampling resolutions.
@@
The different time windows affect the SDR for the SG1 (non-solar) dataset are shown in Figure 26.
@@
The histogram and Gaussian approximation of the change in voltage magnitude for Node 15 of the SG2 solar dataset is shown in Figure 32. This node is consistently predicted to have an incorrect branch by the algorithm when using the Gaussian MI method for varying data lengths. This may be due to the Gaussian approximation being a poor fit as can be seen visually.
@@
The leaf SDR is analysis is shown for 120 day time windows for the SG1 dataset in Figure 34.
@@ Appendix B Kruskal’s Algorithm
Kruskal’s algorithm seeks to find the minimum weight spanning tree. A tree is a graph with no cycles. In a tree you cannot draw a path using connections between nodes on the graph from one node and back to the same node. Each connection is referred to as a pair of nodes $(x,y)$ in the graph and has an associated numerical value known as the weight $w(x,y)$ [cormen2022introduction]. The minimum weight spanning tree is the graph which connects the nodes so that the sum of the weights of the connections is minimized so long as there are no cycles in the graph.
@@
There are several different implementations of Kruskal’s algorithm, two of which differ in how they store the parent of a node [kershenbaum1972computing]. When a pair of nodes is connected in the algorithm, one node is assigned as the parent of the other. This means one node has a higher ranking than the other. A node with four parents could be referred to as a rank four node. Alternatively, if a node has more than one parent, the path to the lowest ranked node can be compressed where the algorithm only keeps track of the lowest ranked parent for each node.
@@
Matlab has its own min-span-tree function but it requires the bio-informatics toolbox which many labs do not have. In order to increase usage of the grid topology estimation algorithm we wrote our own script to avoid licensing issues. Our implementation performs 10x slower than Matlab’s version as seen in Table 4. Since Kruskal’s algorithm execution time is negligible compared to mutual information computation time (as will be soon shown), minimal effort has been placed to improve our implementation of Kruskal’s algorithm.